# H line shape and radius-luminosity relation in 2.5D FRADO

### v1

@@ 7 Conclusion
Following our previous work on the dynamical character and shape of LIL BLR based on the non-hydrodynamic single-cloud 2.5D FRADO model (Naddaf et al., 2021), we here tested the model by calculation of line profiles for the model-concluded distribution of clouds along their trajectories within LIL BLR, for a relatively large grid of initial conditions. The model well predicts the position of the inner radius of the BLR where clouds start to be launched. Predicted line shape factor is roughly consistent with observational trends, we also obtained interesting results for the virial factor. Line shapes are well modeled for larger black hole masses. R-L relation is well reproduced (both the position and the slope) for small Eddington ratios, in case of high Eddington ratio we could possibly recover the observed trend when postulating some selective shielding of the far side of the cloud distribution.
@@ Acknowledgements.
This project is supported by the University of Liege under Special Funds for Research, IPD-STEMA Program. BC acknowledges the OPUS-LAP/GA ˇCR-LA bilateral project (2021/43/I/ST9/01352/OPUS 22 and GF23-04053L). MHN would like to thank Ashwani Pandey for the permission to use his part of CLOUDY results on dependence of H on photon flux. MLMA acknowledges financial support from Millenium Nucleus NCN2023002 (TITANs) an ANID Millennium Science Initiative (AIM23-0001).References
@@
- Du et al. (2015a) Du, P., Hu, C., Lu, K.-X., et al. 2015a, ApJ, 806, 22
- Du et al. (2015b) Du, P., Hu, C., Lu, K.-X., et al. 2015b, ApJ, 806, 22
- Du & Wang (2019) Du, P. & Wang, J.-M. 2019, ApJ, 886, 42
@@
- Goad & Korista (2014) Goad, M. R. & Korista, K. T. 2014, MNRAS, 444, 43
- Goad & Korista (2015) Goad, M. R. & Korista, K. T. 2015, MNRAS, 453, 3662
@@
- Martínez-Aldama et al. (2019) Martínez-Aldama, M. L., Czerny, B., Kawka, D., et al. 2019, ApJ, 883, 170
- Martínez-Aldama et al. (2018) Martínez-Aldama, M. L., del Olmo, A., Marziani, P., et al. 2018, A&A, 618, A179
@@
- Wang et al. (2014a) Wang, J.-M., Qiu, J., Du, P., & Ho, L. C. 2014a, ApJ, 797, 65
- Wang et al. (2014b) Wang, J.-M., Qiu, J., Du, P., & Ho, L. C. 2014b, ApJ, 797, 65

### v3

@@ 7 Conclusion
Building on our previous work concerning the dynamical properties and structure of the low-ionization line (LIL) broad-line region (BLR) using the non-hydrodynamic single-cloud 2.5D FRADO model (Naddaf et al., 2021), we tested the model by computing line profiles based on the predicted distribution of clouds along their trajectories for a relatively broad grid of initial black hole masses and accretion rates. The model is shown to accurately predict the location of the inner radius of the BLR where clouds are launched. The predicted line shape factor is broadly consistent with observational trends and we derived meaningful trends for the virial factor as a function of the model parameters. The line profiles are well reproduced, especially for larger black hole masses. The R-L relation is well captured, both in terms of slope and normalization, for sources with low Eddington ratios. At high Eddington ratios, the observed R-L trend can be qualitatively recovered if the selective obscuration of the far side of the BLR is assumed.
@@ Acknowledgements.
This project is supported by the University of Liege under Special Funds for Research, IPD-STEMA Program. DH is F.R.S.-FNRS Research Director. BC acknowledges the OPUS-LAP/GA ˇCR-LA bilateral project (2021/43/I/ST9/01352/OPUS 22 and GF23-04053L). MHN would like to thank Ashwani Pandey for the permission to use his part of CLOUDY results on dependence of H on photon flux. MLMA acknowledges financial support from Millenium Nucleus NCN2023002 (TITANs) an ANID Millennium Science Initiative (AIM23-0001).References
- Aitken et al. (2000) Aitken, D. K., Greaves, J., Chrysostomou, A., et al. 2000, ApJ, 534, L173
@@
- Du et al. (2015) Du, P., Hu, C., Lu, K.-X., et al. 2015, ApJ, 806, 22
- Du et al. (2016) Du, P., Lu, K.-X., Zhang, Z.-X., et al. 2016, ApJ, 825, 126
- Du & Wang (2019a) Du, P. & Wang, J.-M. 2019a, ApJ, 886, 42
- Du & Wang (2019b) Du, P. & Wang, J.-M. 2019b, ApJ, 886, 42
@@
- Marrone et al. (2007) Marrone, D. P., Moran, J. M., Zhao, J.-H., & Rao, R. 2007, ApJ, 654, L57
- Mart´ınez-Aldama et al. (2019) Martínez-Aldama, M. L., Czerny, B., Kawka, D., et al. 2019, ApJ, 883, 170
- Mart´ınez-Aldama et al. (2018) Martínez-Aldama, M. L., del Olmo, A., Marziani, P., et al. 2018, A&A, 618, A179
@@
- Panda et al. (2018) Panda, S., Czerny, B., Adhikari, T. P., et al. 2018, ApJ, 866, 115
- Panda et al. (2019) Panda, S., Marziani, P., & Czerny, B. 2019, ApJ, 882, 79
- Panda & Śniegowska (2024) Panda, S. & Śniegowska, M. 2024, ApJS, 272, 13
@@
- Wang et al. (2014) Wang, J.-M., Qiu, J., Du, P., & Ho, L. C. 2014, ApJ, 797, 65
@@
- Woo et al. (2024) Woo, J.-H., Wang, S., Rakshit, S., et al. 2024, ApJ, 962, 67
- Yoon et al. (2020) Yoon, D., Chatterjee, K., Markoff, S. B., et al. 2020, MNRAS, 499, 3178